OpenStudy (anonymous):

so the question is: "suppose y = squrt(2x+1) where "x" and "y" are functions of "f". if dx/dt=3 find dy/dt when x=4" I have no clue how to start this, if you start this, perhaps I can finish it

OpenStudy (anonymous):

\[y=\sqrt{2x+1}\]\[\frac{dy}{dt}=\frac{1}{\sqrt{2x+1}}\frac{dx}{dt}\] by the chain rule

OpenStudy (anonymous):